OpenStudy (anonymous):

log(7)3=K 7^(3y-1)=21 howw

OpenStudy (anonymous):

\[\log _{7}3=k\] \[7^{3y-1}=21\]

OpenStudy (anonymous):

What's the question? Get k and y?

OpenStudy (anonymous):

y in terms of K

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[7^{3y-1}=21\] \[7^{3y-1}=7*3\] \[7^{3y-2}=3\] \[log{7^{3y-2}}=log{3}\] \[(3y-2)log{7}=log{3}\] \[(3y-2)=log{3}/log(7)\] \[3y-2=log_7{3}\] \[3y-2=k\] \[3y=k+2\] \[3y=(k+2)/3\]

OpenStudy (anonymous):

Last one is \[y=(k+2)/3\]